Okay, now ETMX's badness is a show-stopper. I'm not sure why, but after this last lockloss, ETMX won't stay put. Right now (as opposed to earlier tonight) it seems to only be happening when I enable LSC pushing on the SUS. ETMX is happy to sit and stay locked on TEM00 green while I write this entry, but if I go and try to turn on the LSC it'll be wacky again. Daytime work.
Anyhow, this is too bad, since I was feelin' pretty good about transitioning DARM over to AS55.
I had a line on (50 counts at 503.1 Hz pushing differentially on the ETMs), and could clearly see the sign flip happen in normalized AS55Q between arm powers of 4 and 6. The line also told me that I needed a matrix element of negative a few x10^-4 in the AS55Q -> DARM spot. Unfortunately, I was missing a zero (so I was making my matrix element too big by a factor of 10) in my ezcastep line, so both times I tried to transition I lost lock.
So. I think that we should put values of 0.5 into the power normalization for our test case (I was using SRCL_IN1 as my tester) since that's the approximate value that the DCtrans uses, and see what size AS55Q matrix element DARM wants tomorrow (tonight was 1.6-3 x 10^-4, but with 1's in the normalization matrix). I feel positive about us getting over to AS55.
Also, Q is (I assume) going to work some more tomorrow on PRMI->REFL165, and Diego is going to re-test his new IR resonance finding script. Manasa, if you're not swamped with other stuff, can you please see if you can have a look at ETMX? Maybe don't change any settings, but see what things being turned on makes ETMX crazy (if it's still happening in the morning).